OpenStudy (anonymous):

10 distinct computational tasks are to be assigned to 3 processors. The first processor should receive 5 tasks, the second processor should receive 3 tasks, and the third processor should receive 2 tasks. The total number of different assignments that obey the given constraint is....

OpenStudy (anonymous):

is it 2520?, (10 choose 5) times (5choose3) times (2choose2)

OpenStudy (2bornot2b):

\[3!\times (^{10}C_5\times ^5C_3\times ^2C_2)\]

OpenStudy (anonymous):

why multiply by 3 factorial?

OpenStudy (2bornot2b):

I am considering that you are allowed to decide which machine you call 1 and which 2 and which 3

OpenStudy (2bornot2b):

Is it OK?

OpenStudy (anonymous):

i understand what your saying but does the question mean that? i mean your answer would be appropriate if its just a division of tasks among 3 processors

OpenStudy (2bornot2b):

The question doesn't state that out clearly. However, you know which answer is for which question..

OpenStudy (anonymous):

k this is the follow up question

OpenStudy (anonymous):

10 distinct computational tasks are to be assigned to 3 processors. The total number of different assignments is...

OpenStudy (2bornot2b):

\[3^{10}\]

OpenStudy (anonymous):

ok

OpenStudy (2bornot2b):

Any more questions? Or I am out of the way..

OpenStudy (mathmate):

3^10 assumes the processors are distinct. (not explicitly specified in the question). If the processors are not distinct (such as same model, same brand), then it is 3^10/3!.

OpenStudy (anonymous):

that works out to be 9841.5 possibilities